In Part 2 of Lab 5 you will dissolve about 3 g of
Answer #1

P1:

mass of solution

mass of water = 1 g/mL * 50 mL = 50 g

total mass of solution = 4.72+50 = 54.72 g

P2.

Assume that the specific heat of the mix is equal to that of water, i.e. 4,184 J/gK

P3.

Absolute value of chang ein temperature:

Tf-Ti = 28.3-25.0 = 3.3 °C

P4.

Total heat involved

Q = m*C*(Tf-Ti)

Q = 54.72*4.184*3.3 = 755.529 J